# DCT-MARL: A Dynamic Communication Topology Based MARL Algorithm for Platoon Control

### v1

@@ V-B Evaluation Metrics
To comprehensively evaluate the performance of the proposed platoon control strategy, this paper adopted two categories of metrics: platoon stability and driving comfort. Platoon stability reflects the system’s ability to suppress inter-vehicle disturbances and avoid amplification of oscillations downstream. Following[21], we define the stability index $S_{i}$ for each vehicle as the time-averaged weighted squared deviation in headway and relative velocity:
@@
$G_{i}^{\prime}=1-\frac{1}{N_{s}}\sum_{t=1}^{N_{s}}\left(\frac{acc_{i}(t)-acc_{i}(t-1)}{jerk_{\text{max}}\cdot\Delta t}\right)^{2},$ (28)
where $jerk_{\text{max}}$ represents the maximum allowable change rate of acceleration used for normalization, which is set to 1.5 m/s3 according to SAE comfort thresholds [39]. Higher values of $G_{i}^{\prime}$ correspond to smoother acceleration transitions and thus improved driving comfort.
@@ VI-A Control Performance against Existing Platoon Control Strategies
To verify the effectiveness and robustness of the proposed strategy, this paper conducts a comprehensive comparison between the proposed DCT-MARL method and other state-of-the-art baselines: the traditional optimization-based Dynamic Weights Optimization based Control (DWOC) strategy [20], which considers both communication delay and packet losse by dynamically compensating for front-vehicle information through link-quality-aware weighting; the DRL-based $D^{3}$PG method [16] proposed a self-adaptive topology control method that significantly improves resilience under fluctuating link conditions. To ensure a fair comparison, the state space setting of $D^{3}$PG is aligned with our proposed method, including the incorporation of delay compensation mechanism.
Fig. 2 illustrates the temporal responses of vehicle velocity and acceleration under an oscillatory driving scenario. Across all methods, a noticeable attenuation of acceleration fluctuations is observed along the platoon, indicating partial suppression of error propagation. However, notable discrepancies persist in the response to abrupt velocity variations and in the smooth attenuation of acceleration during convergence.
(a) DWOC Based Platoon Control
(b) $D^{3}PG$ Based Platoon Control
(c) DCT-MARL Based Platoon Control
@@
The DWOC strategy (Fig. 2 (a)) achieves control stability by adjusting the weights assigned to front vehicle information. This approach provides a certain degree of robustness against delay and packet loss. However, sharp acceleration changes in acceleration are observed between 15s and 25s, indicating challenges in maintaining smooth transitions during sudden maneuvers of the lead vehicle. In contrast, $D^{3}PG$ (Fig. 2 (b)) improves inter-agent coordination by selecting communication neighbors using attention-based mechanisms. Nonetheless, due to its single-agent DRL design lacking full MARL coordination, downstream vehicles show delayed reactions to upstream disturbances. The proposed DCT-MARL method (Fig. 2 (c)) consistently delivers the most stable and coordinated control responses, featuring minimal velocity fluctuations and enhanced inter-vehicle coherence, thereby demonstrating superior robustness under non-ideal communication conditions.
As shown in Table II, DWOC and $D^{3}$PG maintains consistent performance across the platoon; however, DWOC disturbances tend to amplify toward the rear of the platoon, indicating limited suppression capability. Although $D^{3}$PG alleviates this issue via communication-aware strategies, intermittent oscillations still occur under uncertain communication conditions. In contrast, DCT-MARL achieves the lowest string stability metrics across all vehicles, particularly in $CAV_{3}$ to $CAV_{5}$, demonstrating its superior ability to suppress disturbance propagation. Moreover, DCT-MARL yields the highest overall driving comfort scores, maintaining low acceleration variance even for the last vehicle in the platoon, which is typically most susceptible to disturbance, thereby ensuring an improved passenger experience.
@@
To illustrate this more clearly, Fig. 3 presents the performance comparison of various control strategies. The DCT-MARL curve maintains a relatively flat comfort trajectory, while all other baselines show progressively deteriorating performance over time. This highlights DCT-MARL’s robustness and resilience under dynamic and uncertain communication conditions. By jointly considering topology adaptation, delay compensation, and cooperative policy learning, the DCT-MARL method achieves superior multi-objective optimization, balancing safety, comfort, and stability more effectively than DWOC, and $D^{3}PG$. In summary, the DCT-MARL model has been proven to be effective in improving road traffic capacity.
Figure 3: Quantified Metrics of Dfferent Models.
@@ VI-B Control Performance against Existing MARL Algorithms
To further assess the advantages of the proposed DCT-MARL framework, we compare it against two representative MARL algorithms: Delay-Compensated MARL (DEC-MARL)[25], which explicitly addresses communication delay by extending the state space to incorporate historical control outputs under a fixed communication topology. Learning-based MARL (LEB-MARL)[26], which introduces a framework to learn communication priors through causal inference and regularizes agent policies to enhance message utilization.
@@
Table III compares the control performance of the three methods in terms of string stability and driving comfort for each vehicle within the platoon. While DEC-MARL enhances system stability to a certain extent under non-ideal communication conditions, the reliance on a fixed communication topology and absence of dynamic adjustment mechanisms cause significant performance degradation under high packet loss rates and dynamic communication structures. This results in greater spacing fluctuations among following vehicles and a noticeable decline in driving comfort. LEB-MARL incorporates agent correlation and demonstrates certain advantages in control accuracy. However, it fails to account for the stochastic nature of communication links, limiting robustness under packet loss and topological disturbances and leading to more pronounced oscillations in string stability. In contrast, DCT-MARL consistently delivers more robust and stable control outcomes across all vehicles in the platoon.
@@
Besides training performance, the runtime efficiency of the proposed DCT-MARL is also assessed during the testing phase. On a local testing platform equipped with a 2.0 GHz quad-core Intel Core i5 processor and 16 GB of 3733 MHz LPDDR4X memory, executing a complete 600-step simulation required approximately 3.7 seconds. This corresponds to an average inference time of around 6 ms per control step, which is significantly below the 100ms control horizon. These results demonstrate that the proposed method not only achieves superior control performance but also possesses high inference efficiency, making it well-suited for real-time applications or deployment scenarios with limited computational resources.
@@ VI-C Eiffectiveness of Dynamic Communictaion Topology
Figure 5: Communication Heatmap of CAV Platoon with Dynamic Topology.
In the heatmap, the color intensity indicates the frequency of message exchanges between agents. The vertical axis represents the message sender, while the horizontal axis corresponds to the receiver. In this visualization, a darker color indicates a higher likelihood of information exchange, reflecting stronger communication links between specific vehicle pairs. This allows us to intuitively assess how communication patterns evolve under the dynamic topology.
Notably, as the leading vehicle operates as a non-controllable preset agent, it is only possible to observe whether other vehicles receive messages from it, without direct access to the messages received by the leader itself. Accordingly, the corresponding cells representing the leader as a receiver are shaded in light gray to indicate the absence of observable data.
It can be observed that while front-vehicle information receives predominant attention in the control process, rear-vehicle data also demonstrates significant utilization. This observation is consistent with the findings of[5], highlighting the latent value of rear-vehicle states in control decision-making. Compared to traditional fixed topologies, the dynamic topology can adaptively adjust communication links in real-time based on control needs and information value. These results demonstrate the meaningfulness and effectiveness of incorporating dynamic communication topology under non-ideal communication conditions.
@@ VII Conclusions and Future Works
This paper addresses connected vehicle platoon control under non-ideal communication conditions characterized by time-varying delays and packet loss. To tackle these challenges, DCT-MARL is proposed. Specifically, a multi-key gated communication mechanism is introduced, which dynamically adjusts the communication topology based on the correlations among agents and their current communication status to effectively mitigate packet loss. Furthermore, a state augmentation scheme is designed to incorporate historical control actions and communication delay information into each agent’s state representation, ensuring the preservation of the Markov property despite delayed observations. This enhancement further improves the robustness of the system in realistic V2V communication environments. Experimental results demonstrate that the proposed approach significantly improves platoon stability and comfort, outperforming existing state-of-the-art baselines across multiple key performance metrics. Future research will further investigate the mechanism of non-ideal communication and explore the joint optimization of communication and control, enabling more adaptive and resource-efficient platoon coordination.

### v2

@@ V-B Evaluation Metrics
To comprehensively evaluate the performance of the proposed platoon control strategy, this paper adopted two categories of metrics: platoon stability and driving comfort. Platoon stability reflects the system’s ability to suppress inter-vehicle disturbances and avoid amplification of oscillations downstream. Following[21], this paper define the stability index $S_{i}$ for each vehicle as the time-averaged weighted squared deviation in headway and relative velocity:
@@
$G_{i}^{\prime}=1-\frac{1}{N_{s}}\sum_{t=1}^{N_{s}}\left(\frac{acc_{i}(t)-acc_{i}(t-1)}{acc_{\text{max}}}\right)^{2},$ (29)
where $acc_{i}(t)$ denotes the acceleration of vehicle $i$ at time step $t$, and $a_{\text{max}}$ is the maximum allowable acceleration change used for normalization. Higher values of $G_{i}^{\prime}$ correspond to smoother acceleration transitions and thus greater driving comfort.
@@ VI-A Control Performance against Existing Platoon Control Strategies
To verify the effectiveness and robustness of the proposed strategy, this paper conducts a comprehensive comparison between the proposed DCT-MARL method and other state-of-the-art methods. These include the traditional optimization-based Dynamic Weights Optimization Control (DWOC) strategy [20], which accounts for both communication delay and packet loss by dynamically compensating front-vehicle information through communication-quality-aware weighting, and the DRL-based $D^{3}$PG method [16], which introduces a self-adaptive topology control approach that significantly enhances resilience under fluctuating communication conditions. To ensure a fair comparison, the state space settings of D3PG are aligned with our proposed method, including the incorporation of a delay compensation mechanism.
Figure 2 illustrates the temporal responses of vehicle velocity and acceleration under an oscillatory driving scenario. Across all methods, a noticeable attenuation of acceleration fluctuations is observed along the platoon, indicating partial suppression of error propagation. However, notable discrepancies persist in the response to abrupt velocity variations and in the smooth attenuation of acceleration during convergence.
(a) DWOC Policy
(b) $D^{3}PG$ Policy
(c) DCT-MARL Policy
@@
The DWOC strategy (Fig.2 (a)) achieves control stability by adaptively re-weighting front-vehicle information, offering a degree of robustness to delay and packet loss. However, noticeable acceleration spikes between 15s and 25s reveal challenges in ensuring smooth transitions during sudden lead vehicle maneuvers. In contrast, $D^{3}PG$ (Fig.2(b)) improves inter-agent coordination by selecting communication neighbors using attention-based mechanisms. Nonetheless, due to its single-agent DRL design lacking full MARL coordination, downstream vehicles show delayed reactions to upstream disturbances. The proposed DCT-MARL method (Fig.2(c)) consistently delivers the most stable and coordinated control responses, featuring minimal velocity fluctuations and enhanced inter-vehicle coherence, thereby demonstrating superior robustness under non-ideal communication conditions.
Table II summarizes quantitative results. Both DWOC and $D^{3}PG$ maintain reasonable performance across the platoon, yet DWOC exhibits disturbance amplification toward the rear of the platoon, indicating limited suppression capability. Although $D^{3}PG$ alleviates this issue via communication-aware strategies, intermittent oscillations still occur under uncertain communication conditions. In contrast, DCT-MARL achieves the lowest string stability metrics across all vehicles, particularly in $CAV_{3}$ to $CAV_{5}$, demonstrating its superior ability to suppress disturbance propagation. Moreover, DCT-MARL yields the highest overall driving comfort scores, maintaining low acceleration variance even for the last vehicle in the platoon, which is typically most susceptible to disturbance, thereby ensuring an improved passenger experience.
@@
To illustrate this more clearly, Fig.3 presents the performance comparison of various control strategies. The DCT-MARL curve maintains a relatively flat comfort trajectory, while all other baselines show progressively deteriorating performance over time. This highlights DCT-MARL’s robustness and resilience under dynamic and uncertain communication conditions. By jointly considering topology adaptation, delay compensation, and cooperative policy learning, the DCT-MARL method achieves superior multi-objective optimization, balancing safety, comfort, and stability more effectively than DWOC, and $D^{3}PG$.
Figure 3: Quantified Metrics of Different Models.
In summary, the proposed DCT-MARL framework demonstrates significant advantages over existing platoon control methods by effectively mitigating the adverse effects of communication delay and packet loss. It consistently enhances string stability and driving comfort throughout the platoon, especially for vehicles more vulnerable to disturbances.
@@ VI-B Control Performance against Existing MARL Algorithms
To further assess the advantages of the proposed DCT-MARL framework, this paper compare it against two representative MARL algorithms: Delay-Compensated MARL (DEC-MARL) [25], which explicitly addresses communication delay by extending the state space to incorporate historical control outputs under a fixed communication topology. Learning-based MARL (LEB-MARL) [26], which leverages causal inference to infer communication topologies and regularizes agent policies to enhance performance in multi-agent cooperative scenarios.
@@
Table III compares the control performance of the three methods in terms of string stability and driving comfort for each vehicle within the platoon. While DEC-MARL enhances system stability to a certain extent under non-ideal communication conditions, the reliance on a fixed communication topology and absence of dynamic adjustment mechanisms cause significant performance degradation under high packet loss rates and dynamic communication structures. This results in greater spacing fluctuations among following vehicles and a noticeable decline in driving comfort. LEB-MARL incorporates agent correlation and demonstrates certain advantages in control accuracy. However, it fails to account for the time-varying nature of communication quality, limiting robustness under packet loss and topological disturbances and leading to more pronounced oscillations in string stability. In contrast, DCT-MARL consistently delivers more robust and stable control outcomes across all CAVs in the platoon.
@@
Beyond training performance, the runtime efficiency of the proposed DCT-MARL framework is evaluated during testing. On a local platform equipped with a 2.0 GHz quad-core Intel Core i5 processor and 16 GB of 3733 MHz LPDDR4X memory, executing a 600-step simulation required approximately 3.7 seconds, indicating that each decision step takes approximately 6 ms. Given the 100 ms sampling period of vehicle control systems, this inference time is well within the control horizon. These results demonstrate that DCT-MARL not only achieves superior control performance but also exhibits high inference efficiency, rendering it suitable for realtime deployment in computationally constrained environments.
@@ VI-C Eiffectiveness of Dynamic Communictaion Topology
Figure 5: Communication Heatmap of Connected Platoon with Dynamic Topology.
In the heatmap, the color intensity indicates the frequency of message exchanges between agents. The vertical axis represents the message sender, while the horizontal axis corresponds to the receiver. In this visualization, a darker color indicates a higher likelihood of information exchange, reflecting stronger communication between specific vehicle pairs. This allows us to intuitively assess how communication patterns evolve under the dynamic topology.
Notably, as the leading vehicle operates as a noncontrollable preset agent, it is only possible to observe whether other vehicles receive messages from it, without direct access to the messages received by the leader itself. Accordingly, the corresponding cells representing the leader as a receiver are shaded in light gray to indicate the absence of observable data.
It can be observed that while front-vehicle information receives predominant attention in the control process, rear vehicle data also demonstrates significant utilization. This observation is consistent with the findings of [5], highlighting the latent value of rear vehicle states in control decision making. Compared to traditional fixed topologies, the dynamic topology can adaptively adjust communication agents in real time based on control needs and information value. These results demonstrate the meaningfulness and effectiveness of incorporating dynamic communication topology under non-ideal communication conditions.
@@ VII Conclusions and Future Works
This paper proposes a DCT-MARL framework to address connected vehicle platoon control under non-ideal V2V communications characterized by time-varying delay and packet loss. Specifically, a multi-key gated communication mechanism is introduced, which dynamically adjusts the communication topology based on the correlations among vehicles and their current communication status to effectively mitigate packet loss. Furthermore, a state augmentation scheme is designed to incorporate historical control actions and communication delay information into each agent’s state space, ensuring the preservation of the Markov property despite delayed observations. This enhancement further improves the robustness of the system in realistic V2V communication environment. Experimental results demonstrate that the proposed approach significantly improves platoon stability and driving comfort of vehicle platoon compared to existing advanced methods. Future research will further investigate the mechanism of non-ideal communication and explore the joint optimization of communication and control, enabling more adaptive and resource-efficient platoon coordination.